OpenStudy (anonymous):

the numerator of a fraction is 2 less than its denominator. When both numerator and denominator are increased by 3, the fraction is increased by3/20. Find the original fraction.

OpenStudy (badhi):

If the denominator is x, "the numerator of a fraction is 2 less than its denominator" => \[\frac{x-2}{x}\] "When both numerator and denominator are increased by 3"=> \[\frac{x+1}{x+3}\] "the fraction is increased by3/20"=> \[\frac{x-2}{x}+\frac 3{20}\] \[\frac{x+1}{x+3}=\frac{x-2}{x}+\frac 3{20}\] Solve this